OpenStudy (anonymous):

Solve for the indicated variable. Show all work. A = LW ; for L (I'm sure this is easy but I really cannot understand how to figure this out.)

OpenStudy (anonymous):

A= LW this is a actually very simple when you look at it. You want to get L by itself but L is being multiplied by W. So in order to get rid of W we're going to divide it on both sides. A/W= L/W the W attached to the L cancels out so then we're left with A/W=L